# Learning Theory for Kernel Bilevel Optimization

### v1

@@ 4 Generalization Bounds for (KBO)
In this section, we present the main result of the present work: a maximal inequality controlling how both value function $\mathcal{F}$ and its gradient $\nabla\mathcal{F}$ are well approximated by their empirical counter-parts uniformly over a compact subset $\Omega$ of $\mathbb{R}^{d}$. In Section 4.1, we state the main result and discuss its implications. Then we present the general proof strategy, in Section 5, and discuss possible alternative approaches.
@@ 4.1 Maximal Inequalities for (KBO)
The following theorem provides finite sample bounds on the uniform approximation errors on the objective and its gradient in expectation over both inner and upper-level samples.
@@
Fix any compact subset $\Omega$ of $\mathbb{R}^{d}$. Under Assumptions (B), (D), (A) and (C), the following maximal inequalities hold:
$\mathbb{E}\left[\sup_{\omega\in\Omega}\left|\mathcal{F}(\omega)-\widehat{\mathcal{F}}(\omega)\right|\right]$ $\leq C{\left(\frac{1}{\sqrt{m}}+\frac{1}{\sqrt{n}}\right)},$
@@
$\mathbb{E}\left[\sup_{\omega\in\Omega}\left\|\nabla\mathcal{F}(\omega)-\widehat{\nabla\mathcal{F}}(\omega)\right\|\right]$ $\leq C{\left(\frac{1}{\sqrt{m}}+\frac{1}{\sqrt{n}}\right)},$
@@
where the expectation is taken over the finite samples and $C$ is a constant that depends only $\Omega$, dimension $d$, regularization parameter $\lambda$, $\kappa$ and local upper-bounds on $\ell_{in}$, $\ell_{out}$ and their partial derivatives over suitable compact set.
@@
Theorem 4.1 states that the estimation error can be decomposed into two contributions each resulting from finite sample approximation of $L_{in}$ and $L_{out}$ with a parametric rate of $\frac{1}{\sqrt{n}}$ and $\frac{1}{\sqrt{m}}$ up to a constant factor $C$. We provide a detailed expression for the constant in Theorem D.7 of Appendix D. The restriction to compact subsets $\Omega$ instead of the whole space $\mathbb{R}^{d}$ allows controlling the complexity of some function classes indexed by the parameter $\omega$. Without additional assumptions on the objectives, we obtain a constant $C$ that grow with the diameter of the subset $\Omega$.
@@
To illustrate the implications of Theorem 4.1, we provide two convergence results for bilevel gradient methods in Section 4.2 and defer the discussion on the general proof strategy to Section 5 with a full proof provided in Appendix D.
@@ 4.2 Applications to Empirical Bilevel Gradient Methods
A typical strategy to solve (KBO) is to obtain empirical samples and solve ($\widehat{\text{KBO}}$) using a bilevel optimization algorithm. Note that this is a finite-dimensional problem, and the lower-level is typically strongly convex. Our results allow to provide statistical guarantees for such approaches. We start with the simplest possible gradient algorithm for the unconstrained problem, $\mathcal{C}=\mathbb{R}^{d}$, given $\eta>0$:
$\omega_{t+1}$ $=\omega_{t}-\eta\nabla\widehat{\mathcal{F}}(\omega_{t})$ $t\geq 0.$ (6)
The algorithm requires access to the (almost surely) strongly convex inner-level solution and its derivative which can be obtained using implicit differentiation.
@@
Consider Assumptions (B), (D), (A) and (C) and a fixed $\lambda>0$. Assume furthermore that $\mathbf{K}$ in ($\widehat{\text{KBO}}$) is almost surely definite and that there is $c>0$ such that $\inf_{\omega,v,y}\ell_{out}(\omega,v,y)-c\|\omega\|^{2}>-\infty$. Fix $\omega_{0}\in\mathbb{R}^{d}$ and assume that $\omega_{t}$ is given by (6) for all $t\geq 0$. Then there are $\bar{\eta}>0$ and a constant $\bar{c}>0$ such that for any $0<\eta<\bar{\eta}$, and for any $t>0$,
$\mathbb{E}\Big{[}\min_{i=0,\ldots,t}\left\|\nabla\mathcal{F}(\omega_{i})\right\|\Big{]}$ $\leq\bar{c}\left(\frac{1}{\sqrt{m}}+\frac{1}{\sqrt{n}}+\frac{1}{\sqrt{t+1}}\right),$
@@
$\mathbb{E}\Big{[}\underset{i\to\infty}{\lim\sup}\left\|\nabla\mathcal{F}(\omega_{i})\right\|\Big{]}$ $\leq\bar{c}\left(\frac{1}{\sqrt{m}}+\frac{1}{\sqrt{n}}\right).$
Proof.
Consider that $\inf_{\omega,v,y}\ell_{out}(\omega,v,y)-c\|\omega\|^{2}\geq 0$, which entails $\ell_{out}(\omega,v,y)\geq c\|\omega\|^{2}$ for all $v,y$. Using Proposition C.1 and setting $B=\sup_{y\in\mathcal{Y}}|\partial_{v}\ell_{in}(\omega_{0},0,y)|$, we have almost surely
^$\widehat{\mathcal{F}}(\omega_{0})\leq\max_{|v|\leq B\kappa/\lambda,y\in\mathcal{Y}}\ell_{out}(\omega_{0},v,y)\coloneqq\bar{\ell}.$
[…]
  • ^$\widehat{\mathcal{F}}(\omega_{t})\leq\widehat{\mathcal{F}}(\omega_{0})$ and $\omega_{t}\in\Omega$ for all $t\geq 0$.
  • ^$\nabla\widehat{\mathcal{F}}(\omega_{t})\to 0$ as $t\to\infty$.
The corollary then follows by combining Proposition 3.1 and the uniform bound in Theorem 4.1. ∎
Remark 4.3.
The additional assumption on $\ell_{out}$ is a device to ensure a priori almost sure boundedness of the sequence. It is rather mild as it can be enforced by a small perturbation of the form $(\omega,v,y)\mapsto\ell_{out}(\omega,v,y)+c\|\omega\|^{2}$ assuming $\ell_{out}\geq 0$, which is typical in applications. Any other device ensuring a priori boundedness could be considered. The assumption on $\operatorname{\mathbf{K}}$ is satisfied almost surely for most kernels.
Now, considering the constrained problem and assuming $\mathcal{C}$ is convex compact, the projected gradient descent initialized at $\omega_{0}\in\mathcal{C}$, given $\eta>0$, iterates the following recursion:
$\omega_{t+1}=\Pi_{\mathcal{C}}(\omega_{t}-\eta\nabla\widehat{\mathcal{F}}(\omega_{t}))$
for all $t\geq 0$, where $\Pi_{\mathcal{C}}$ denotes the orthogonal projection on $\mathcal{C}$. The algorithmic requirements are the same as the gradient algorithm, with the addition of the orthogonal projection, which is a cheap operation for basic sets such as balls. For constrained optimization, the optimality condition should take the constraints into account. We consider the gradient mapping, see (Beck, 2017, Section 10.3),
$\widehat{G}_{\eta}$ $\colon\omega\mapsto\frac{1}{\eta}\left(\omega-\Pi_{\mathcal{C}}(\omega-\eta\nabla\widehat{\mathcal{F}}(\omega))\right),$
$G_{\eta}$ $\colon\omega\mapsto\frac{1}{\eta}\left(\omega-\Pi_{\mathcal{C}}(\omega-\eta\nabla\mathcal{F}(\omega))\right).$
This captures stationarity for the recursion, and any local minimum of $\mathcal{F}$ on $\mathcal{C}$ satisfies $G_{\eta}=0$ for all $\eta>0$.
Corollary 4.4 (Generalization for bilevel projected gradient descent).
Consider Assumptions (B), (D), (A) and (C) and a fixed $\lambda>0$. Assume furthermore that $\mathbf{K}$ in ($\widehat{\text{KBO}}$) is almost surely definite and $\mathcal{C}$ is convex compact. Assume that $\omega_{t+1}=\omega_{t}-\eta\nabla\widehat{\mathcal{F}}(\omega_{t})$ for all $t\geq 0$. Then there are $\bar{\eta}>0$ and a constant $\bar{c}>0$ such that for any $0<\eta<\bar{\eta}$, and for any $t>0$,
$\mathbb{E}\Big{[}\min_{i=0,\ldots,t}\left\|G_{\eta}(\omega_{i})\right\|\Big{]}$ $\leq\bar{c}\left(\frac{1}{\sqrt{m}}+\frac{1}{\sqrt{n}}+\frac{1}{\sqrt{t+1}}\right),$
$\mathbb{E}\Big{[}\underset{i\to\infty}{\lim\sup}\left\|G_{\eta}(\omega_{i})\right\|\Big{]}$ $\leq\bar{c}\left(\frac{1}{\sqrt{m}}+\frac{1}{\sqrt{n}}\right).$
Proof.
We choose $\Omega=\mathcal{C}$. All iterates obviously remain in $\Omega$. Similarly as in the proof of Corollary 4.2, $\nabla\widehat{\mathcal{F}}$ is Lipschitz and $\mathcal{F}$ is bounded on $\Omega$ with deterministic constants. It then follows using classical analysis of nonconvex projected gradient algorithm (see, e.g., (Beck, 2017, Theorem 10.15)), that for small $\eta$, almost surely $\min_{i=0,\ldots,t}\left\|\widehat{G}_{\eta}(\omega_{i})\right\|\leq\bar{c}/\sqrt{t+1}$ for a determinstic $\bar{c}>0$, and that $\widehat{G}_{\eta}(\omega_{i})\to 0$ as $i\to\infty$. Using the fact that the orthogonal projection is $1$-Lipschitz, (see, e.g., (Beck, 2017, Theorem 6.42)), we have for all $\omega\in\Omega$
The result follows by combining Proposition 3.1 and the uniform bound in Theorem 4.1. ∎
5 General Proof Strategy for Theorem 4.1
The main strategy behind the proof of Theorem 4.1 in Appendix D consists in three steps: (step 1) obtaining a point-wise error decomposition of the errors into manageable error terms that holds almost surely for any $\omega\in\Omega$, then applying maximal inequalities to suitable empirical processes (step 2) and some degenerate $U$-processes (step 3) to control each of these terms. The final error bounds are obtained by combining all these bounds as shown in Section D.3.
Step 1: Point-wise error decomposition. A main challenge in controlling the errors in Theorem 4.1 is the non-linear dependence of both estimators ^$\widehat{\mathcal{F}}(\omega)$ and ^$\widehat{\nabla\mathcal{F}}(\omega)$ on the empirical distributions, as they are obtained via a plug-in procedure. We address this challenge by breaking down the error into components based on the discrepancies between expected values and their empirical counterparts of individual point-wise losses and their derivatives, evaluated at the optimal solution $h_{\omega}^{\star}$. Specifically, we denote by $\delta_{\omega}^{out}$ and $\delta_{\omega}^{in}$ the errors on the objectives defined as:
$\delta_{\omega}^{out}$ $\coloneqq{\left\lvert L_{out}(\omega,h^{\star}_{\omega})-\widehat{L}_{out}(\omega,h^{\star}_{\omega})\right\rvert},$
$\delta_{\omega}^{in}$ $\coloneqq{\left\lvert L_{in}(\omega,h^{\star}_{\omega})-\widehat{L}_{in}(\omega,h^{\star}_{\omega})\right\rvert}.$
Additionally, we quantify the errors between the partial derivatives of these objectives and their empirical counterparts. To simplify our proof outline, we slightly abuse notation by denoting $\partial_{h}\delta_{\omega}^{out}$, $\partial_{h}\delta_{\omega}^{in}$, $\partial_{\omega}\delta_{\omega}^{out}$, $\partial_{\omega,h}^{2}\delta_{\omega}^{in}$ and $\partial_{h}^{2}\delta_{\omega}^{in}$ to refer to these errors in terms of partial derivatives. For instance, $\partial_{h}\delta_{\omega}^{out}$ is defined as ${\left\lVert\partial_{h}L_{out}(\omega,h^{\star}_{\omega})-\partial_{h}\widehat{L}_{out}(\omega,h^{\star}_{\omega})\right\rVert}_{\mathcal{H}}$, with similar definitions for the other error terms that can be found in Section D.1. The next proposition formalizes the above discussions.
Proposition 5.1 (Approximation bounds).
Let $\Omega$ be an arbitrary compact subset of $\mathbb{R}^{d}$. Under Assumptions (B), (D), (A) and (C), the following holds almost surely, for any $\omega\in\Omega$:
${\left\lvert\mathcal{F}(\omega)-\widehat{\mathcal{F}}(\omega)\right\rvert}$ $\leq C{\left(\delta_{\omega}^{out}+\partial_{h}\delta_{\omega}^{in}\right)},$
${\left\lVert\nabla\mathcal{F}(\omega)-\widehat{\nabla\mathcal{F}}(\omega)\right\rVert}$ $\leq C(\partial_{\omega}\delta_{\omega}^{out}+\partial_{h}\delta_{\omega}^{out}+\partial_{h}^{2}\delta_{\omega}^{in}$
$\qquad+\partial_{\omega,h}^{2}\delta_{\omega}^{in}+\partial_{h}\delta_{\omega}^{in}),$
where $C$ is a positive constant that depends only $\Omega$, dimension $d$, regularization parameter $\lambda$, $\kappa$ and local upper-bounds on $\ell_{in}$, $\ell_{out}$ and their partial derivatives over suitable compact set.
A more detailed version of this proposition, including the exact constants, along with its proof are provided in Proposition D.4 of Section D.1. The error terms arising in the above decomposition are more amenable to a statistical analysis using empirical process theory as we discuss next.
Step 2: Maximal inequalities for empirical processes. Some of the error terms, namely $\delta_{\omega}^{out}$ and $\partial_{\omega}\delta_{\omega}^{out}$, can be controlled directly using empirical process theory. For instance, $\delta_{\omega}^{out}$ is associated to the family of random functions $\sqrt{m}{\left(L_{out}(\omega,h_{\omega}^{\star})-\widehat{L}_{out}(\omega,h_{\omega}^{\star})\right)}_{\omega\in\Omega}$, which defines an empirical process, a scaled and centered empirical average of real-valued functions indexed by the parameter $\omega$. Thus, provided that suitable estimates of the complexity of the class are available (as measured by its packing number in Proposition E.1 of Appendix E), which is easy to obtain in our setting, we show in Proposition D.5 of Section D.2 that a maximal inequality of the form below follows from classical results on empirical processes:
$\mathbb{E}_{\mathbb{Q}}{\left[\sup_{\omega\in\Omega}\overbrace{{\left\lvert L_{out}(\omega,h_{\omega}^{\star})-\widehat{L}_{out}(\omega,h_{\omega}^{\star})\right\rvert}}^{\delta_{\omega}^{out}}\right]}\leq\frac{C}{\sqrt{m}}.$
Step 3: Maximal inequalities for degenerate U-processes. Unfortunately, the approach in step 2 cannot be readily used for the remaining error terms involving partial derivatives w.r.t. $h$ ($\partial_{h}\delta_{\omega}^{out}$, $\partial_{h}\delta_{\omega}^{in}$, $\partial_{\omega,h}^{2}\delta_{\omega}^{in}$ and $\partial_{h}^{2}\delta_{\omega}^{in}$). These error terms are associated to processes that are not real-valued anymore but take values in an infinite-dimensional space. While the recent work in (Park & Muandet, 2023) develops an empirical process theory for functions taking values in a vector space, the provided complexity estimates would result in unfavorable dependence on the sample size. Instead, we leverage the structure of the RKHS to control these errors using maximal inequalities for suitable degenerate $U$-processes of order $2$ indexed by the parameter $\omega$ and for which such inequalities were provided in the seminal works of Sherman (1994); Nolan & Pollard (1987). $U$-processes of order 2 are generalization of empirical processes that involve empirical averages of real-valued functions which depend on pairs of samples, instead of a single one as in empirical processes. These functions arise, in our case, when taking the square of the error term $\partial_{h}\delta_{\omega}^{out}$, for instance, and exploiting the reproducing property in the RKHS. This approach, presented in Proposition D.6 of Section D.2, allows us to obtain maximal inequalities of the form:
$\mathbb{E}_{\mathbb{Q}}{\left[\sup_{\omega\in\Omega}\partial_{h}\delta_{\omega}^{out}\right]}\leq\mathbb{E}_{\mathbb{Q}}{\left[\sup_{\omega\in\Omega}{\left(\partial_{h}\delta_{\omega}^{out}\right)}^{2}\right]}^{\frac{1}{2}}\leq\frac{C}{\sqrt{m}}.$
Combining the maximal inequalities from step 2 and 3 with the error decomposition from step 1 allows to obtain the result of Theorem 4.1.
Discussion. Alternative approaches to $U$-processes could be used to derive generalization bounds, although these would result in degraded sample dependence. Specifically, one could employ a variational formulation of the RKHS norm appearing in some of the error terms, such as $\partial_{h}\delta_{\omega}^{out}$, to express them as the error of some real-valued empirical process to which standard results could be applied. However, this comes at the cost of considering processes indexed not only by the finite-dimensional parameter $\omega$, but also by functions in a unit RKHS ball, in contrast with our approach based on $U$-processes indexed by finite-dimensional vectors. Consequently, these families have much larger complexities as measured by their covering/packing numbers (Yang et al., 2020, Lemma D.2), which directly impacts the generalization rate. Our proposed approach bypasses this challenge by using real-valued $U$-processes indexed by finite-dimensional parameters, at the expense of employing a more general empirical process theory for degenerate $U$-processes (Sherman, 1994).

### v3

@@ 4 Generalization bounds for (KBO)
In this section, we present our main result: a maximal inequality that controls how well both $\mathcal{F}$ and $\nabla\mathcal{F}$ are approximated by their empirical counterparts, uniformly over a compact subset $\Omega$ of $\mathbb{R}^{d}$.
@@ 4.1 Maximal inequalities for (KBO)
The following theorem provides finite-sample bounds on the uniform approximation errors on the objective and its gradient in expectation over both inner- and outer-level samples.
@@
Fix any compact subset $\Omega$ of $\mathbb{R}^{d}$. Under Assumptions˜(A), (C), (E), (B) and (D), the following maximal inequalities hold:
@@
$\mathbb{E}\left[\sup_{\omega\in\Omega}\left|\mathcal{F}(\omega)-\widehat{\mathcal{F}}(\omega)\right|\right]\leq C\left(\frac{1}{\sqrt{m}}+\frac{1}{\sqrt{n}}\right),\mathbb{E}\left[\sup_{\omega\in\Omega}\left\|\nabla\mathcal{F}(\omega)-\widehat{\nabla\mathcal{F}}(\omega)\right\|\right]\leq C\left(\frac{1}{\sqrt{m}}+\frac{1}{\sqrt{n}}\right)$
@@
where the expectation is taken over the finite samples, and $C$ is a constant that depends only on $\Omega$, the dimension $d$, the regularization parameter $\lambda$, $\kappa$, and local upper bounds on $\ell_{in}$, $\ell_{out}$, and their partial derivatives over suitable compact sets.
@@
Theorem˜4.1 states that the estimation error can be decomposed into two contributions each resulting from finite-sample approximation of $L_{in}$ and $L_{out}$ with a parametric rate of $\nicefrac{{1}}{{\sqrt{m}}}$ and $\nicefrac{{1}}{{\sqrt{n}}}$ up to a constant factor $C$. We provide a detailed expression for the constant in Theorem˜E.7 of Appendix˜E. The restriction to a compact subset $\Omega$ instead of the whole space $\mathbb{R}^{d}$ allows controlling the complexity of some function classes indexed by the parameter $\omega$. Without further assumptions on the objectives, we obtain a constant $C$ that can grow with the diameter of the subset $\Omega$.
Role of $\lambda$. The regularization parameter $\lambda$ simultaneously controls the strong convexity of the inner objective (see Proposition˜B.3 of Section˜B.1), the boundedness and Lipschitz continuity of the inner solutions (see Section˜D.1), the smoothness of the outer objective (see Proposition˜D.4 of Section˜D.2), the modulus of continuity of $L_{in}$, $\widehat{L}_{in}$, $L_{out}$, $\widehat{L}_{out}$ and their partial derivatives (see Section˜E.1), and maximal inequalities for certain processes (see Section˜E.2). Larger values of $\lambda$ yield smoother problems that are faster to optimize with larger step sizes, but introduce a larger statistical bias, while smaller values of $\lambda$ reduce bias but make optimization and generalization more delicate, with the error tending to $\infty$ as $\lambda\to 0$. Selecting $\lambda$ therefore involves a trade-off between bias (regularized vs un-regularized problems), variance (finite sample vs population, as done in our study), and optimization efficiency (step size). Under our assumptions, the dependence of the constants on $\lambda$ cannot be quantified, precluding generalization guarantees as $\lambda\to 0$. Stronger assumptions are required to obtain a quantitative control on $\lambda$.
Probabilistic and variance bounds. The most difficult quantity to control is the expectation of the maximal differences, which we have established. Once this expectation is bounded, a high-probability bound on the maximal differences can be derived via Markov’s inequality. Moreover, since these differences are bounded (see Proposition˜E.4), we can also bound their variance. Indeed, let $Z\in[0,z]$ be a random variable representing the maximal difference between $\mathcal{F}$ or $\nabla\mathcal{F}$ and their respective plug-in estimators. Then, $\operatorname{\text{Var}}(Z/z)\leq\mathbb{E}[(Z/z)^{2}]\leq\mathbb{E}[Z/z]$, which implies that $\operatorname{\text{Var}}(Z)\leq z\mathbb{E}[Z]$.
@@
We outline the general proof strategy for Theorem˜4.1 in the following section, with a full proof provided in Appendix˜E.
4.2 General proof strategy for Theorem˜4.1
The main strategy behind the proof of Theorem˜4.1 in Appendix˜E consists of three steps: (step 1) obtaining a point-wise error decomposition of the errors into manageable error terms that holds almost surely for any $\omega\in\Omega$, then applying maximal inequalities to suitable empirical processes (step 2) and some degenerate $U$-processes (step 3) to control each of these terms. The final error bounds are obtained by combining all these bounds as shown in Section˜E.3.
Step 1: point-wise error decomposition. A main challenge in controlling the errors in Theorem˜4.1 is the non-linear dependence of both estimators $\widehat{\mathcal{F}}(\omega)$ and $\widehat{\nabla\mathcal{F}}(\omega)$ on the empirical distributions, as they are obtained via a plug-in procedure. We address this by breaking down the errors into components based on the discrepancies between expected values and their empirical counterparts of individual point-wise losses and their derivatives, all evaluated at the optimal solution $h_{\omega}^{\star}$. Specifically, we denote by $\delta_{\omega}^{out}$ and $\delta_{\omega}^{in}$ the errors on the objectives defined as $\delta_{\omega}^{out}\coloneqq|L_{out}(\omega,h^{\star}_{\omega})-\widehat{L}_{out}(\omega,h^{\star}_{\omega})|$ and $\delta_{\omega}^{in}\coloneqq|L_{in}(\omega,h^{\star}_{\omega})-\widehat{L}_{in}(\omega,h^{\star}_{\omega})|$. Moreover, we quantify the errors between the partial derivatives of these objectives and their empirical counterparts. To simplify our proof outline, we slightly abuse notation by denoting $\partial_{h}\delta_{\omega}^{out}$, $\partial_{h}\delta_{\omega}^{in}$, $\partial_{\omega}\delta_{\omega}^{out}$, $\partial_{\omega,h}^{2}\delta_{\omega}^{in}$, and $\partial_{h}^{2}\delta_{\omega}^{in}$ to refer to these errors in terms of partial derivatives. For instance, $\partial_{h}\delta_{\omega}^{out}$ is defined as $\|\partial_{h}L_{out}(\omega,h^{\star}_{\omega})-\partial_{h}\widehat{L}_{out}(\omega,h^{\star}_{\omega})\|_{\mathcal{H}}$, with similar definitions for the other terms (see Section˜E.1). We get $|\mathcal{F}(\omega)-\widehat{\mathcal{F}}(\omega)|\leq C(\delta_{\omega}^{out}+\partial_{h}\delta_{\omega}^{in})$ and $\|\nabla\mathcal{F}(\omega)-\widehat{\nabla\mathcal{F}}(\omega)\|\leq C(\partial_{\omega}\delta_{\omega}^{out}+\partial_{h}\delta_{\omega}^{out}+\partial_{h}^{2}\delta_{\omega}^{in}+\partial_{\omega,h}^{2}\delta_{\omega}^{in}+\partial_{h}\delta_{\omega}^{in})$. Proposition˜E.4 formalizes this step and includes the exact constants. The error terms in both decompositions are amenable to a statistical analysis using empirical process theory as we discuss next.
Step 2: maximal inequalities for empirical processes. Some of the error terms, namely $\delta_{\omega}^{out}$ and $\partial_{\omega}\delta_{\omega}^{out}$, can be controlled directly using empirical process theory. For example, $\delta_{\omega}^{out}$ is associated to the family of random functions $\sqrt{m}(L_{out}(\omega,h_{\omega}^{\star})-\widehat{L}_{out}(\omega,h_{\omega}^{\star}))_{\omega\in\Omega}$, which defines an empirical process, a scaled and centered empirical average of real-valued functions indexed by the parameter $\omega$. Thus, provided that suitable estimates of the class complexity are available (as measured by its packing number in Proposition˜F.1 of Appendix˜F), which are easy to obtain in our setting, we show in Proposition˜E.5 of Section˜E.2 that a maximal inequality of the following form follows from classical results on empirical processes: $\mathbb{E}_{\mathbb{Q}}\left[\sup_{\omega\in\Omega}\delta_{\omega}^{out}\right]\leq\nicefrac{{C}}{{\sqrt{m}}}$.
Step 3: maximal inequalities for degenerate U-processes. Step 2 cannot be readily applied to the remaining terms involving partial derivatives w.r.t. $h$ $(\partial_{h}\delta_{\omega}^{out},\partial_{h}\delta_{\omega}^{in},\partial_{\omega,h}^{2}\delta_{\omega}^{in},\partial_{h}^{2}\delta_{\omega}^{in})\eqqcolon\mathbf{D}_{h}$. These are associated to processes that are not real-valued anymore, but take values in an infinite-dimensional space. In fact, one could apply step 2 to get an error per dimension, but then summing the errors yields a divergent sum. While the recent work in [park2023towards] develops an empirical process theory for functions taking values in a vector space, the provided complexity estimates would result in an unfavorable dependence on the sample size. Instead, we leverage the structure of the RKHS to control these errors using maximal inequalities for suitable degenerate $U$-processes of order $2$ indexed by the parameter $\omega$ and for which such inequalities were provided in the seminal works of sherman1994maximal, nolan1987u. $U$-processes of order 2 are generalization of empirical processes and involve empirical averages of real-valued functions which depend on pairs of samples, instead of a single one as in empirical processes. In our case, these functions arise when taking the square of any term in $\mathbf{D}_{h}$ and exploiting the reproducing property of the RKHS. This approach, presented in Proposition˜E.6 of Section˜E.2, allows us to obtain maximal inequalities for the terms in $\mathbf{D}_{h}$. For example, it is of the following form for $\partial_{h}\delta_{\omega}^{out}$: $\mathbb{E}_{\mathbb{Q}}\left[\sup_{\omega\in\Omega}\partial_{h}\delta_{\omega}^{out}\right]\leq\nicefrac{{C}}{{\sqrt{m}}}$. Combining the maximal inequalities from steps 2 and 3 with the error decomposition from step 1 allows to obtain the result of Theorem˜4.1.
Discussion. Alternative approaches to $U$-processes could be used to derive generalization bounds, although these would result in a degraded sample dependence. Specifically, one could employ a variational formulation of the RKHS norm appearing in some of the error terms, such as $\partial_{h}\delta_{\omega}^{out}$, to express them as the error of some real-valued empirical process to which standard results could be applied. However, this comes at the cost of considering processes indexed not only by the finite-dimensional parameter $\omega$, but also by functions in the unit RKHS ball. As a result, these families have much larger complexities as measured by their covering/packing numbers [yang2020function, Lemma D.2], which directly impacts the generalization rate. In contrast, our proposed approach bypasses this challenge by using real-valued $U$-processes indexed by finite-dimensional parameters, at the expense of employing a more general empirical process theory for degenerate $U$-processes [sherman1994maximal].
To illustrate the implications of Theorem˜4.1, we next provide convergence results for bilevel gradient methods.
@@ 4.3 Applications to empirical bilevel gradient methods
A typical strategy to solve (KBO) is to obtain empirical samples and apply a bilevel optimization algorithm to ($\widehat{\text{KBO}}$), for which our results offer statistical guarantees. Below, we present the generalization error for bilevel gradient descent. Generalization results for the projected bilevel gradient descent are directed to Section˜E.4.
Bilevel gradient descent. It is the simplest gradient-based method for solving the unconstrained (KBO) problem, i.e., when $\mathcal{C}=\mathbb{R}^{d}$. It performs the update $\omega_{t+1}=\omega_{t}-\eta\nabla\widehat{\mathcal{F}}(\omega_{t})$ for all $t\geq 0$, where $\eta>0$ is the step size. The algorithm requires access to the strongly convex inner-level solution and its derivative, which can be obtained using implicit differentiation.
@@
Consider Assumptions˜(A), (C), (E), (B) and (D) and fix $\lambda>0$. Assume further that $\mathbf{K}$ in ($\widehat{\text{KBO}}$) is almost surely definite, and that there exists $c>0$ such that $\inf_{\omega,v,y}\ell_{out}(\omega,v,y)-c\|\omega\|^{2}>$-$\infty$. Fix $\omega_{0}\in\mathbb{R}^{d}$ and let $\omega_{t+1}=\omega_{t}-\eta\nabla\widehat{\mathcal{F}}(\omega_{t})$ for all $t\geq 0$, where $\eta>0$ is the step size. Then, there exist constants $\bar{\eta}>0$ and $\bar{c}>0$ such that for any $0<\eta<\bar{\eta}$ any $t>0$, the following holds:
@@
$\mathbb{E}\Big[\min_{i=0,\ldots,t}\left\|\nabla\mathcal{F}(\omega_{i})\right\|\Big]\leq\bar{c}\left(\frac{1}{\sqrt{m}}+\frac{1}{\sqrt{n}}+\frac{1}{\sqrt{t+1}}\right),\mathbb{E}\Big[\underset{i\to\infty}{\lim\sup}\left\|\nabla\mathcal{F}(\omega_{i})\right\|\Big]\leq\bar{c}\left(\frac{1}{\sqrt{m}}+\frac{1}{\sqrt{n}}\right)$
The additional assumption on $\ell_{out}$ serves as a device to ensure the almost sure boundedness of the sequence a priori. It is rather mild, as it can be enforced by a small perturbation of the form $(\omega,v,y)\mapsto\ell_{out}(\omega,v,y)+c\|\omega\|^{2}$, assuming $\ell_{out}\geq 0$, which is typical in applications. Any other device that ensures a priori boundedness could also be considered. The assumption on $\operatorname{\mathbf{K}}$ is satisfied almost surely for most commonly used kernels. The proof of Corollary˜4.2 follows from Theorem˜4.1 and can be found in Section˜E.4. Corollary˜4.2 also highlights a key algorithmic insight: the convergence of the bilevel method requires striking a balance between data availability (sample sizes $n$ and $m$) and computational budget (number of gradient steps). This trade-off arises because the total convergence error combines a statistical component, due to approximating the population gradient from finite samples, and an optimization component, due to performing only a limited number of (projected) gradient steps. Ensuring the right balance when designing practical algorithms prevents either insufficient data or limited computation from dominating the overall error.
5 Numerical experiments
Setup. To empirically validate our theoretical results, we consider the instrumental variable regression problem discussed in Section˜2.2, in which we assume a linear dependence on $\omega$ for the function $f_{\omega}$ of the form $f_{\omega}(t)=\omega^{\top}\phi(t)$, where $\phi(t)\in\mathbb{R}^{d}$ denotes the feature map. We chose this particular problem because it allows us to derive closed-form expressions of the exact value function and its gradient, as well as their plug-in estimators, which are detailed in Section˜I.2. We use the Gaussian kernel and follow the experimental setup of singh2019kernel, generating synthetic data that remain fixed across all runs. We vary $n$ between $100$ and 5,0005{,}000, setting $m=n$. The case $n\neq m$ is analyzed separately in Section˜I.5. For the instrumental variable $x$, we consider two distributions: a $p$-dimensional standard Gaussian and a $p$-dimensional Student’s $t$-distribution with degrees of freedom $\nu\in\{2.1,2.5,2.9\}$. Further details on the experimental setup are provided in Section˜I.4. We optimize the outer loss in ($\widehat{\text{KBO}}$) using gradient descent, where the step size is selected using backtracking line search and $\omega_{0}$ is randomly drawn from $\mathcal{U}(0,1)^{d}$. The stopping criterion is when $\|\nabla\widehat{\mathcal{F}}(\omega_{i})\|\leq 10^{-5}$, where $i$ is the iteration index. Our code is available at https://github.com/fareselkhoury/KBO.
Scalable approximations for $\mathcal{F}$ and $\nabla\mathcal{F}$. Since the expressions of $\mathcal{F}$ and $\nabla\mathcal{F}$ involve expectations, they are intractable to compute exactly. We approximate them accurately using their plug-in estimators $\widehat{\mathcal{F}}$ and $\widehat{\nabla\mathcal{F}}$, derived in Section˜I.2, and evaluate them using a large number of samples. To make this computation scalable, we approximate the kernel using a random Fourier features approximation [rahimi2007random], as detailed in Section˜I.3. Specifically, we use 1,000,0001{,}000{,}000 samples and 26,00026{,}000 random features, and handle memory constraints via a block decomposition strategy with a block size of 1,0001{,}000.
Results. The plots in Figure˜2 show the generalization behavior as a function of the number of inner samples $n$. (a) and (b) display the generalization error at initialization for the value function and the gradient, respectively. (c) presents the generalization bound for the gradient norm at the final iteration, while (d) shows the bound for the minimum gradient norm across all iterations. These results align with our theoretical findings, as all curves closely follow the expected theoretical slope. Additionally, Figure˜3 of Section˜I.5 shows that balanced sample sizes lead to improved optimization behavior.
Figure 2: Illustration of gradient descent on ($\widehat{\text{KBO}}$) for the instrumental variable regression task using synthetic data. The plots are averaged over 50 runs and displayed on a log-log scale. The line represents the mean across all runs, and the shaded region indicates the 95% confidence interval.